MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2009/11/13
OBSERVER: Hayashi
CONDITION: clear
COMMENT: At start CCD temperature was about -50. But it was -60 while taking dark.
Compressor gas pressure: 1.90 MPa
Expected Observation Time: START 22:11 END 04:51
Go to the Seeing Log
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 23015-23017 | 3 | 7 | 420 |
LMC | 4544-4555 | 12 | 69 | 6900 |
SMC | 888-890 | 3 | 5 | 660 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 18 | 81 | 7980 (02h13m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 0 | 0 | 0 | 0 | 1 | 1 | 0 | 1 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
3 | 3 | 3 | 3 | 3 | 3 | 12 | 12 | 12 | 3 | 3 | 3 |
Field13 | Field14 | Field15 | Field16 |
3 | 3 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
3 | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
22:11 | observing | clear |
23:00 | observing | clear |
00:00 | observing | clear |
01:00 | observing | clear |
02:00 | observing | clear |
03:00 | observing | clear |
04:00 | observing | clear |
04:51 | observing | dawn |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day